OpenStudy (anonymous):

There is a 40 percent chance of rain today and a 50 percent chance of rain tomorrow. If the two events are independent, what is the probability that it will rain today or tomorrow?

OpenStudy (anonymous):

when dealing with independent events, the probability of one event occurring does not affect the probability of another event occurring

OpenStudy (anonymous):

Here you have \[P(rain today) = 40%\] while \[P(rain tomorrow) = 50%\]

OpenStudy (anonymous):

the probability of two independent events occurring is the product of their probabilities: \[P(raintoday)*P(raintomorrow) = P(rain today and tomorrow)\]

OpenStudy (anonymous):

that would b 2000

OpenStudy (anonymous):

2000%?

OpenStudy (amistre64):

doesnt the term "or" indicate addition of probabilities?

OpenStudy (anonymous):

err

OpenStudy (anonymous):

haha, yes!

OpenStudy (anonymous):

sorry, I was giving you the answer for how to determine if both events occur

OpenStudy (anonymous):

so, as amistre pointed out "or" implies addition while "and" implies multiplication

OpenStudy (anonymous):

so, you actually want to \[P(raintoday) + P(raintomorrow) = P(rain today and tomorrow)\]

OpenStudy (anonymous):

I believe

OpenStudy (amistre64):

that looks better, and since the events are independant there is no need to subtract out the P(AnB), maybe

OpenStudy (anonymous):

right, so, \[P(A or B) = P(A) + P(B) - P(AandB)\] and independence implies that \[P(AandB)=0\]